# Quantitative Imaging of 55Co and 18F-Labeled Tracers in a Single “Multiplexed” PET Imaging Session

### v1

@@ 2.3 Animal Study
Six-week-old female Balb/c mice, sourced from Jackson Labs (Sacramento, CA), were used in the animal experiments. A syngeneic mouse colon cancer line, CT26, was implanted subcutaneously (1 million cells in 0.1 mL of saline) in the right shoulders of the mice 7 days prior to the study. When the tumors reached 100–200 cubic millimeters, probes were prepared and injected retro-orbitally. Co-55 was complexed to a DOTA-aPD-L1 monoclonal antibody conjugate (150 $\mu$Ci/mg specific activity) that binds to the PD-1 antigen on T-cells. However, it appears that there may have been some labeling of an unconjugated DOTA species in the reaction mixture leading to rapid renal clearance and high signal in the kidneys. Despite this limitation, the experiments were still informative for evaluating the proposed mPET signal unmixing strategies. Due to the rapid clearance of the Co-55-labeled antibody, FDG (80 $\mu$Ci in 0.1 mL) was injected 60 minutes before imaging, and the Co-55-labeled antibody (25 $\mu$Ci in 0.1 mL) was injected 15 minutes before imaging. Single-tracer and dual-tracer mice were prepared, and PET scans were conducted as described above. At the end of the experiments, the mice were sacrificed.
@@ 4.2 Animal Scans
The Co-55 image in the dual-tracer mouse (Fig. 4) is noisy due to lower statistics of triple coincidences. The accumulation of FDG in the body is consistent with previous literature [15], showing high uptake in the heart, brain, and bladder. The accumulation of free Co-55 in the kidneys is not described in the literature thus leading to our assumption of a Co-DOTA species [16, 17]. However, despite the unexpected biodistribution of the Co-55-labeled antibody, this study still enables us to evaluate the unmixing methodology for the dual-tracer study. The unmixed dual-tracer image (Fig. 4) clearly shows separated tracer distributions that are similar to those observed in the overlaid single-tracer mouse distributions (Fig. 4). The small differences between the two sets of distributions can be attributed to the use of three different mice. The ability to recover similar tracer distributions in the dual-tracer mouse as in the single-tracer mice indicates that the unmixing methodology for simultaneous Co-55 and FDG imaging was successful.

### v2

@@ 2.3 Animal Study
Six-week-old female Balb/c mice, sourced from Jackson Labs (Sacramento, CA), were used in the animal experiments. A syngeneic mouse colon cancer line, CT26, was implanted subcutaneously (1 million cells in 0.1 mL of saline) in the right shoulders of the mice 7 days prior to the study. When the tumors reached 100–200 cubic millimeters, probes were prepared and injected retro-orbitally. Co-55 was complexed to a DOTA-aPD-L1 monoclonal antibody conjugate (150 $\mu$Ci/mg specific activity) that binds to the PD-L1 antigen on tumor cells. However, it appears that there may have been some unstable Co-55 antibody complex in the final product, leading to rapid clearance and high signal in the kidneys. Despite this limitation, the experiments were still informative for evaluating the proposed mPET signal unmixing strategies. Due to the rapid clearance of the Co-55-labeled antibody, FDG (80 $\mu$Ci in 0.1 mL) was injected 60 minutes before imaging, and the Co-55-labeled antibody (25 $\mu$Ci in 0.1 mL) was injected 15 minutes before imaging. Single-tracer and dual-tracer mice were prepared, and PET scans were conducted as described above. At the end of the experiments, the mice were sacrificed.
@@ 4.2 Animal Scans
The Co-55 image in the dual-tracer mouse (Fig. 4) is noisy due to lower statistics of triple coincidences. The accumulation of FDG in the body is consistent with previous literature [15], showing high uptake in the heart, brain, and bladder. The accumulation of free Co-55 in the kidneys is not described in the literature thus leading to our assumption of unstable Co-55 antibody complex in the final product [16, 17]. However, despite the unexpected biodistribution of the Co-55-labeled antibody, this study still enables us to evaluate the unmixing methodology for the dual-tracer study. The unmixed dual-tracer image (Fig. 4) clearly shows separated tracer distributions that are similar to those observed in the overlaid single-tracer mouse distributions (Fig. 4). The small differences between the two sets of distributions can be attributed to the use of three different mice. The ability to recover similar tracer distributions in the dual-tracer mouse as in the single-tracer mice indicates that the unmixing methodology for simultaneous Co-55 and FDG imaging was successful.